OpenStudy (zackattack350):

What is the equation in standard form of the line which passes through (−2, 6) and has a slope of −1?

OpenStudy (zackattack350):

x + y = −4

OpenStudy (zackattack350):

x + y = 4

OpenStudy (zackattack350):

x − y = −4

OpenStudy (zackattack350):

x-y=4

OpenStudy (zackattack350):

I need help learning how to solve this

OpenStudy (texaschic101):

standard form : Ax + By = C In y = mx + b form, the slope is in the m position slope(m) = -1 (-2,6)...x = -2 and y = 6 now we sub 6 = -1(-2) + b 6 = 2 + b 6 - 2 = b 4 = b so your equation is : y = -x + 4..but we need it in Ax + By = C y = -x + 4 --- add x to both sides x + y = 4 <==== your standard form line

OpenStudy (zackattack350):

Thank you so much for explaining it out :) @texaschic101

OpenStudy (texaschic101):

no problem...you got to learn someway :)